# RoDE: Linear Rectified Mixture of Diverse Experts for Food Large Multi-Modal Models

### v1

@@ Introduction
Low-Rank Adaptation (LoRA).
Fully tuning large models can be resource-intensive. Parameter-Efficient Fine-Tuning (PEFT) (Ding et al. 2023) introduces additional adapters to effectively customize large models for downstream tasks with minimal resource overhead. Let $W$ represent the weight of a linear layer $L$ in the large-scale model, which is initially set as $W_{0}$ after pre-training. In PEFT, $W_{0}$ remains frozen in order to preserve the acquired world knowledge. To facilitate adaptation to downstream tasks, a learnable branch linear adapter, denoted as $\Delta W$, is introduced to modify the initial frozen linear weights $W_{0}$. Consequently, the output of the adapted linear layer $L$ can be expressed as $W=W_{0}+\Delta W$. LoRA (Hu et al. 2021) further decomposes $\Delta W$ into two matrices, $A$ and $B$, where the connection rank is considerably smaller than that of $W_{0}$. LoRA allows the model to adapt to downstream tasks while reducing the number of training parameters.
Mixture of Experts (MoE).
The Mixture of Experts (MoE) (Dou et al. 2023) paradigm of LoRA is proposed to adapt large models to multiple downstream tasks by employing multiple LoRAs in the MLP layers. Each LoRA module can be considered as an expert, wherein all experts receive the same input and combine their outputs to improve overall performance. Let $R=\left\{A_{i},B_{i}\right\}_{i=0}^{N}$ denote a series of LoRA modules, where $N$ signifies the number of LoRAs. We use $h(\cdot)$ to denote the linear router, which takes the dimension of the input feature $x$ and outputs the allocation vector of $R$. The output of layer $L$ incorporating the MoE module can be represented as:
@@
Linear Rectified Mixture of Diverse Experts
@@
Diverse Capability Experts
@@
To mitigate the issue of GPU memory constraints, we conceptualize the experts as fine-grained skill modules. The key idea is that a task may activate a combination of these modules, and these modules can be shared across various tasks. This modular design intuitively leads us to develop LoRA experts with different capabilities tailored to tasks of varying complexity. Drawing inspiration from Low-Rank Adaptation (Hu et al. 2021), which suggests that a low-rank adapter may be sufficient for certain tasks, we create LoRAs with varying ranks. Consequently, the resulting skill space comprises both high-rank and low-rank LoRA experts, providing greater flexibility and efficiency in addressing a diverse set of tasks.
@@
Figure 5: Comparison between LLaVA-MoLE (Chen, Jie, and Ma 2024) and our proposed RoDE. In our RoDE framework, experts are utilized across all tasks. For individual tasks, expert activations are optimized to achieve sparsity.
Linear Rectified Router
The router consolidates the contributions of each expert based on the demands of the specific task. Previous research (Zhou et al. 2022) has shown that the use of sparse mixtures of LoRA experts outperforms dense experts in the context of large language models. LLaVA-MoLE (Chen, Jie, and Ma 2024) presents a top-1 selection strategy, which ensures the sparsity of expert selection in LMM. While some methodologies in the Natural Language Processing (NLP) field adopt a ’soft’ approach to combine expert outputs—for instance, (Dou et al. 2023) utilizes softmax and (Ponti et al. 2023) employs Gumbel softmax—these techniques are not as sparse and can be challenging to optimize.
In contrast, our methodology adopts a novel approach by utilizing the Rectified Linear Unit (ReLU) (Nair and Hinton 2010) to rectify the output of the routers, thereby encouraging the sparse learning of LoRA expert activations. A visual illustration of our routing strategy is shown in Figure 5. Leveraging the intrinsic properties of the Rectified Linear Unit (ReLU), our approach benefits from a simplified optimization landscape and fosters sparsity within the network, which enables our model to boost both efficiency and efficacy.
@@
IoU ()($\uparrow$)( ↑ ) SacreBLEU ()($\uparrow$)( ↑ ) Rouge-L ()($\uparrow$)( ↑ ) Mass ()($\downarrow$)( ↓ ) Energy ()($\downarrow$)( ↓ ) Fat ()($\downarrow$)( ↓ ) Proteins ()($\downarrow$)( ↓ ) Carb ()($\downarrow$)( ↓ ) Avg.()($\downarrow$)( ↓ )
FoodLMM (Yin et al. 2023) 22.59 7.59 36.44 48.49 47.72 61.49 56.61 51.25 53.11
FoodLMM+MoE 24.54 8.7 38.5 49.23 48.24 60.59 58.37 51.87 53.66
RoDE (ours) 26.86 9.66 40.11 48.18 46.96 60.96 58.16 48.65 52.58
Table 2: Comparative results for ingredient recognition, recipe generation and nutrition estimation on the Uni-Food Dataset. The FoodLMM+MoE is derived by integrating a Mixture of Experts (MoE) paradigm into the FoodLMM model.
Optimization Objects and Inference
Following previous LMM-based methods (Liu et al. 2023a), the input images are transformed into image tokens and concatenated with text tokens to send to LLM. The training process adheres to the autoregressive model, predicting the next token based on the input tokens. We employ the standard cross-entropy loss as the optimization objective to train our model.
@@
Nutrition5k Mass ()($\downarrow$)( ↓ ) Cal ()($\downarrow$)( ↓ ) Fat ()($\downarrow$)( ↓ ) Proteins ()($\downarrow$)( ↓ ) Carb ()($\downarrow$)( ↓ ) Avg. ()($\downarrow$)( ↓ )
FoodLMM(Yin et al. 2023) 38.96 54.04 69.84 54.88 51.3 53.8
FoodLMM+MoE 39.31 53.6 67.14 52.99 49.59 52.53
RoDE (ours) 38.45 52.47 67.1 53.92 47.84 51.96
Table 3: Nutrition estimation results on Nutrtion5k dataset. The evaluation metric employed is the percentage Mean Absolute Error (pMAE).
Experiment Setup
Our approach commenced by utilizing the pre-trained weights of the LLaVA-Lightning-7B-v1-1 (Liu et al. 2023a) model, subsequently applying SFT on the Uni-Food and Nutrition5k datasets (Thames et al. 2021). In our RoDE model, we employ a heterogeneous array of experts with LoRA ranks of [2, 4, 8, 16]. Our primary baseline is FoodLMM (Yin et al. 2023), which to our knowledge is the only versatile LMM specialized in food-related tasks. Additionally, we compare with FoodLMM+MoE, which substitutes the plain LoRA module with a Mixture of Experts (MoE) module. Each MoE module consists of four LoRAs with a rank of 8. The routing mechanism for the MoE module is a linear layer followed by a softmax operation, as detailed in Section Mixture of Experts (MoE).. Importantly, for simplicity, we omit the specialized token tags and additional heads used by FoodLMM specifically for nutrition estimation, and instead directly convert the output tokens from the LMM into text to generate nutritional predictions. For the standardization of the ingredient vocabulary, we adopt the same clustering protocol used for ingredients in InverseCooking (Salvador et al. 2019).
Tasks and Evaluation Metrics. The tasks and corresponding metrics in our experiments are as follows. Ingredient Recognition: The objective here is to identify and list the ingredients present in a dish shown in an image. We assess performance using Intersection over Union (IoU) as the metric. Recipe Generation: The goal of this task is to generate cooking instructions for the dish captured in an image. We measure performance using the commonly used text generation metrics in NLP, SacreBLEU (Post 2018) and Rouge-L (Lin 2004). Nutrition Estimation: This task involves estimating a range of nutritional parameters from a dish’s image, including the total food mass, total energy, total fat, total carbohydrates, and total protein content. Energy is measured in kilocalories, while the remaining parameters are quantified in grams. We employ the mean absolute error as a percent of the respective mean for that field (pMAE) (Thames et al. 2021) for the metric of nutrition estimation.
Figure 6: Sparsity Heatmap Comparison. This heatmap visualizes the disparity in task allocation sparsity between routers using the softmax strategy and those employing our proposed linear rectification method. The y-axis denotes the middle transformer blocks, while the x-axis represents the output of the routers.
@@
Table 4: Ablation Study on MoE. The integration of MoE can enhance the LMM, not solely due to the introduction of additional LoRA parameters
@@
Table 5: Comparison of Three Routing Strategies. Our proposed Linear Rectified Router demonstrates superior performance on both homogeneous and heterogeneous combinations of experts.
Implementation Details. In this study, we utilize LLaVA-7B (Liu et al. 2023a) as the base LMM, wherein a CLIP ViT-L model serves as the vision encoder, and Llama 7B (Touvron et al. 2023) serves as the Language Model. For the CLIP ViT-L encoder, the input image resolution is set at 336x336, which is then divided into 14 patches. Subsequently, a two-layer MLP is used to transform these image patches into 576 tokens. For efficiency, the LoRA module is added to the query and key projection layer. Unless specifically stated, the rank of LoRA is set to 8. Hyperparameter $\alpha$ is set to 16, and the LoRA dropout rate is set to 0.05.
During the training process of all our experiments, the weights of vision encoder and Llama are kept constant, with only the added LoRA modules being trainable. We employ the AdamW optimizer (Loshchilov and Hutter 2017) and use the WarmupDecayLR learning rate scheduler. The initial learning rate is set to 0.0003, weight decay is set to 0, and we perform 100 warm-up iterations. In the training stage, we use a batch size of 4, and gradient updates are computed every 10 batches. The training process is distributed across 4 RTX4090 GPUs. The training epoch is set to 1.
Performance Comparison
We perform a comparative analysis between our model and two versions of FoodLMM: a base version equipped with plane LoRA, and an enhanced version, FoodLMM+MoE, equipped with four 8-rank LoRA and a linear router with softmax. These methods are evaluated across three tasks: ingredient recognition, recipe generation, and nutrition estimation. The results are summarized in Table 2. The table clearly shows that the MoE design enhances the performance of FoodLMM. By adopting the traditional MoE paradigm, i.e., FoodLMM+MoE, there is an 8.6% improvement in IoU for ingredients recognition, for recipe generation, there is a 14.6% increase in SacreBLEU and a 1.1% gain in Rouge-L scores. However, there is a slight reduction of 1% in nutritional performance. Overall, the comparison demonstrates the advantages of incorporating MoE.
@@
Additionally, we conduct further experiments on Nutrition5k  (Thames et al. 2021) dataset, adhering to the methodology established by FoodLMM, where only image information was utilized for training (without depth information). The outcomes, as presented in Table 3, unequivocally demonstrate that our methods consistently surpass both FoodLMM and traditional MoE variations in performance.
@@
Table 6: Ablation Study on the Rank Settings of Experts within MoE. The heterogeneous combination of experts yields superior performance while utilizing a reduced parameter count.
Ablation study
@@
MoE vs. Larger LoRA rank
@@
As illustrated in Table 4, we first compare the standard LoRA (Var1) with its MoE variant (Var3). The results indicate a superior performance of the MoE design. To ascertain whether this performance improvement can be attributed to an increase in the number of trainable LoRA parameters, we augment the LoRA rank of Var1 to 32 to obtain Var2. A comparison between Var2 and Var4 reveals the superiority of the MoE design, even though these two variants have an equivalent number of trainable LoRA parameters. This finding suggests that the successful integration of MoE is not simply a result of expanding the number of trainable parameters.
Routing Strategy
@@
Then we examine the effects of the routing strategy. Our experiment includes three routing strategies: 1) similar to (Chen, Jie, and Ma 2024), the top-1 expert is selected, referred to as Top-1; 2) following the approach in (Dou et al. 2023), we employ a softmax function to normalize the router outputs, which is denoted as Softmax; 3) we introduce our novel linear rectified router, which we denote with the abbreviation LR. The results are summarized in Table 5 and visualized in Figure 7. As observed, while the Top-1 allocation MoE outperforms the standard LoRA (Var1), it is surpassed by the ”soft” allocation strategies (i.e.Var2 and Var3). This indicates the efficiency of employing an ensemble of experts to address a single task. Moreover, the comparison between the LR and Softmax routers (i.e. Var2 and Var3, Var4 and Var5) suggests that the LR routing strategy is superior. These results further underscore the effectiveness of our linear rectification routing strategy.
@@
In addition, we visualize the heatmaps of router outputs for both the Softmax and LR routing strategies across several middle transformer blocks. Figure 6 illustrates this comparative analysis, clearly indicating that our proposed Linear Rectified Router achieves a higher degree of sparsity in task allocation.
Rank configuration of LoRA

### v3

@@ IV-A Preliminary
IV-A2 Low-Rank Adaptation (LoRA).
Fully tuning large models can be resource-intensive. Parameter-Efficient Fine-Tuning (PEFT) [42] introduces additional adapters to effectively customize large models for downstream tasks with minimal resource overhead. Let $W$ represent the weight of a linear layer $L$ in the large-scale model, which is initially set as $W_{0}$ after pre-training. In PEFT, $W_{0}$ remains frozen in order to preserve the acquired world knowledge. To facilitate adaptation to downstream tasks, a learnable branch linear adapter, denoted as $\Delta W$, is introduced to modify the initial frozen linear weights $W_{0}$. Consequently, the output of the adapted linear layer $L$ can be expressed as $W=W_{0}+\Delta W$. LoRA [30] further decomposes $\Delta W$ into two matrices, $A$ and $B$, where the connection rank is considerably smaller than that of $W_{0}$. LoRA allows the model to adapt to downstream tasks while reducing the number of training parameters.
IV-A3 Mixture of Experts (MoE).
The Mixture of Experts (MoE) [26] paradigm of LoRA is proposed to adapt large models to multiple downstream tasks by employing multiple LoRAs in the MLP layers. Each LoRA module can be considered as an expert, wherein all experts receive the same input and combine their outputs to improve overall performance. Let $R=\left\{A_{i},B_{i}\right\}_{i=0}^{N}$ denote a series of LoRA modules, where $N$ signifies the number of LoRAs. We use $h(\cdot)$ to denote the linear router, which takes the dimension of the input feature $x$ and outputs the allocation vector of $R$. The output of layer $L$ incorporating the MoE module can be represented as:
@@
IV-B Linear Rectified Mixture of Diverse Experts
@@
IV-B1 Diverse Capability Experts
@@
To mitigate the issue of GPU memory constraints, we conceptualize the experts as fine-grained skill modules. The key idea is that a task may activate a combination of these modules, and these modules can be shared across various tasks. This modular design intuitively leads us to develop LoRA experts with different capabilities tailored to tasks of varying complexity. Drawing inspiration from Low-Rank Adaptation [30], which suggests that a low-rank adapter may be sufficient for certain tasks, we create LoRAs with varying ranks. Consequently, the resulting skill space comprises both high-rank and low-rank LoRA experts, providing greater flexibility and efficiency in addressing a diverse set of tasks.
@@
Figure 6: Comparison between LLaVA-MoLE [23] and our proposed RoDE. In our RoDE framework, experts are utilized across all tasks. For individual tasks, expert activations are optimized to achieve sparsity.
IV-B2 Linear Rectified Router
The router consolidates the contributions of each expert based on the demands of the specific task. Previous research [22] has shown that the use of sparse mixtures of LoRA experts outperforms dense experts in the context of large language models. LLaVA-MoLE [23] presents a top-1 selection strategy, which ensures the sparsity of expert selection in LMM. While some methodologies in the Natural Language Processing (NLP) field adopt a “soft” approach to combine expert outputs—for instance, [26] utilizes softmax and [28] employs Gumbel softmax—these techniques are not as sparse and can be challenging to optimize.
In contrast, our methodology adopts a novel approach by utilizing the Rectified Linear Unit (ReLU) [43] to rectify the output of the routers, thereby encouraging the sparse learning of LoRA expert activations. A visual illustration of our routing strategy is shown in Figure 6. Leveraging the intrinsic properties of the Rectified Linear Unit (ReLU), our approach benefits from a simplified optimization landscape and fosters sparsity within the network, which enables our model to boost both efficiency and efficacy.
@@ IV-B2 Linear Rectified Router
TABLE III: Comparative results for ingredient recognition, recipe generation and nutrition estimation on the Uni-Food Dataset. The FoodLMM+MoE is derived by integrating a Mixture of Experts (MoE) paradigm into the FoodLMM model.
@@
IoU ()($\uparrow$)( ↑ ) F1 ()($\uparrow$)( ↑ ) SacreBLEU ()($\uparrow$)( ↑ ) Rouge-L ()($\uparrow$)( ↑ ) Mass ()($\downarrow$)( ↓ ) Energy ()($\downarrow$)( ↓ ) Fat ()($\downarrow$)( ↓ ) Proteins ()($\downarrow$)( ↓ ) Carb ()($\downarrow$)( ↓ ) Avg.()($\downarrow$)( ↓ )
QwenVL [44] 20.01 30.51 7.97 39.19 49.99 48.12 65.19 62.95 49.41 55.13
FoodLMM [18] 22.59 32.98 7.59 36.44 48.49 47.72 61.49 56.61 51.25 53.11
FoodLMM+MoE 24.54 35.21 8.7 38.5 49.23 48.24 60.59 58.37 51.87 53.66
RoDE (ours) 26.86 37.67 9.66 40.11 48.18 46.96 60.96 58.16 48.65 52.58
TABLE IV: Results on Food-101, VireoFood-172 and Nutrition 5k. We utilize Food-101 for the classification task, VireoFood-172 for the ingredient recognition task, and Nutrition 5k for nutrition estimation task. The term “FoodLMM-s1” refers to the version of FoodLMM that has been fine-tuned on the target dataset. Similarly, “RoDE-ft” denotes the fine-tuned version of RoDE.
Method Food 101 VireoFood-172 Nutrition 5k
Acc ()($\uparrow$)( ↑ ) IoU ()($\uparrow$)( ↑ ) F1 ()($\uparrow$)( ↑ ) Mass ()($\downarrow$)( ↓ ) Energy ()($\downarrow$)( ↓ ) Fat ()($\downarrow$)( ↓ ) Proteins ()($\downarrow$)( ↓ ) Carb ()($\downarrow$)( ↓ ) Avg.()($\downarrow$)( ↓ )
PRENet [35] 91.13 - - - - - - - -
CACLNet [10] - - 65.71 - - - - - -
FoodLMM-s1 [18] 93.93 56.94 68.97 38.96 54.04 69.84 54.88 51.3 53.8
FoodLMM-s1+MoE 94.1 61.38 69.36 39.31 53.6 67.14 52.99 49.59 52.53
RoDE-ft (ours) 94.02 65.95 73.54 38.45 52.47 67.1 53.92 47.84 51.96
IV-C Optimization Objects and Inference
Following previous LMM-based methods [17], the input images are transformed into image tokens and concatenated with text tokens to send to LLM. The training process adheres to the autoregressive model, predicting the next token based on the input tokens. We employ the standard cross-entropy loss as the optimization objective to train our model.
@@ V Experiment
V-A Experiment Setup
Our approach commenced by utilizing the pre-trained weights of the LLaVA-Lightning-7B-v1-1 [17] model, subsequently applying SFT on the Uni-Food and Nutrition5k datasets [21]. In our RoDE model, we employ a heterogeneous array of experts with LoRA ranks of [2, 4, 8, 16]. Our primary baseline is FoodLMM [18], which to our knowledge is the only versatile LMM specialized in food-related tasks. Additionally, we compare with FoodLMM+MoE, which substitutes the plain LoRA module with a Mixture of Experts (MoE) module. Each MoE module consists of four LoRAs with a rank of 8. The routing mechanism for the MoE module is a linear layer followed by a softmax operation, as detailed in Section IV-A3. Furthermore, we explore the capabilities of QwenVL [44], a LMM that is not ideally suited for the food domain. We fine-tune QwenVL on the UniFood dataset using pretrained weights. Importantly, for simplicity, we omit the specialized token tags and additional heads used by FoodLMM specifically for nutrition estimation, and instead directly convert the output tokens from the LMM into text to generate nutritional predictions. For the standardization of the ingredient vocabulary, we adopt the same clustering protocol used for ingredients in InverseCooking [37].
V-A1 Tasks and Evaluation Metrics
The tasks and corresponding metrics in our experiments are as follows. Ingredient Recognition: The objective here is to identify and list the ingredients present in a dish shown in an image. We assess performance using Intersection over Union (IoU) as the metric. Recipe Generation: The goal of this task is to generate cooking instructions for the dish captured in an image. We measure performance using the commonly used text generation metrics in NLP, SacreBLEU [45] and Rouge-L [46]. Nutrition Estimation: This task involves estimating a range of nutritional parameters from a dish’s image, including the total food mass, total energy, total fat, total carbohydrates, and total protein content. Energy is measured in kilocalories, while the remaining parameters are quantified in grams. We employ the mean absolute error as a percent of the respective mean for that field (pMAE) [21] for the metric of nutrition estimation.
Figure 7: Comparison of Sparsity Heatmaps for Expert Activation. This heatmap visualizes the disparity in task allocation sparsity between routers using the softmax strategy and those employing our proposed linear rectification method. The y-axis denotes the middle transformer blocks, while the x-axis represents the output of the routers.
TABLE V: Ablation Study on MoE. The integration of MoE can enhance the LMM, not solely due to the introduction of additional LoRA parameters
@@ V-A1 Tasks and Evaluation Metrics
TABLE VI: Comparison of Three Routing Strategies. Our proposed Linear Rectified Router demonstrates superior performance on both homogeneous and heterogeneous combinations of experts.
@@
V-A2 Implementation Details
In this study, we utilize LLaVA-7B [17] as the base LMM, wherein a CLIP ViT-L model serves as the vision encoder, and Llama 7B [16] serves as the Language Model. For the CLIP ViT-L encoder, the input image resolution is set at 336x336, which is then divided into 14 patches. Subsequently, a two-layer MLP is used to transform these image patches into 576 tokens. For efficiency, the LoRA module is added to the query and key projection layer. Unless specifically stated, the rank of LoRA is set to 8. Hyperparameter $\alpha$ is set to 16, and the LoRA dropout rate is set to 0.05.
During the training process of all our experiments, the weights of vision encoder and Llama are kept constant, with only the added LoRA modules being trainable. We employ the AdamW optimizer [47] and use the WarmupDecayLR learning rate scheduler. The initial learning rate is set to 0.0003, weight decay is set to 0, and we perform 100 warm-up iterations. In the training stage, we use a batch size of 4, and gradient updates are computed every 10 batches. The training process is distributed across 4 RTX4090 GPUs. The training epoch is set to 1.
V-B Performance Comparison
V-B1 Results on UniFood
We conduct a comparative analysis among our model, QwenVL, and two versions of FoodLMM: a base version equipped with plane LoRA, and an enhanced version, FoodLMM+MoE, equipped with four 8-rank LoRA and a linear router with softmax. These methods are evaluated across three tasks: ingredient recognition, recipe generation, and nutrition estimation. The summarized outcomes in Table III reveal that FoodLMM outperforms QwenVL in ingredient recognition and nutrition estimation tasks but falls short in recipe generation. The integration of the Mixture of Experts (MoE) framework significantly bolsters FoodLMM’s capabilities. Specifically, the FoodLMM+MoE version, which employs the classical MoE approach, achieves an 8.6% enhancement in Intersection over Union (IoU) for ingredient recognition. For recipe generation, it exhibits a notable 14.6% increase in SacreBLEU and a 5.6% improvement in Rouge-L scores. Nonetheless, there is a marginal decrease of 1% in the nutritional estimation performance. Overall, these comparative results underscore the benefits of incorporating the MoE paradigm into FoodLMM.
@@ V-B1 Results on UniFood
V-B2 Results on other datasets
Additionally, we extend our evaluation to specific single tasks, including food classification, ingredient recognition, and nutrition estimation, utilizing distinct datasets for each task: the Food-101 dataset [32] for food classification, the VireoFood-172 dataset [40] for ingredient recognition, and the Nutrition 5k dataset [21] for nutrition estimation. For the Food-101 dataset, our method is benchmarked against state-of-the-art methods including PRENet [35] and FoodLMM [18]. In the case of the VireoFood-172 dataset, we draw comparisons with CACLNet [10] and FoodLMM. Regarding the Nutrition 5k dataset, we adhere to the methodology established by FoodLMM, which relies solely on image information for training, deliberately excluding depth data to maintain consistency with their approach. The results, detailed in Table IV, clearly demonstrate that our methods consistently outperform both traditional methods and FoodLMM across all evaluated tasks. This consistent superiority in performance underscores the effectiveness of our approach in tackling diverse challenges within the food domain.
@@ V-B2 Results on other datasets
TABLE VII: Ablation Study on the Rank Settings of Experts within MoE. The heterogeneous combination of experts yields superior performance while utilizing a reduced parameter count.
@@
V-C Ablation study
@@
V-C1 MoE vs. Larger LoRA rank
@@
As illustrated in Table V, we first compare the standard LoRA (Var1) with its MoE variant (Var3). The results indicate a superior performance of the MoE design. To ascertain whether this performance improvement can be attributed to an increase in the number of trainable LoRA parameters, we augment the LoRA rank of Var1 to 32 to obtain Var2. A comparison between Var2 and Var4 reveals the superiority of the MoE design, even though these two variants have an equivalent number of trainable LoRA parameters. This finding suggests that the successful integration of MoE is not simply a result of expanding the number of trainable parameters.
V-C2 Routing Strategy
@@
Then we examine the effects of the routing strategy. Our experiment includes three routing strategies: 1) similar to [23], the top-1 expert is selected, referred to as Top-1; 2) following the approach in [26], we employ a softmax function to normalize the router outputs, which is denoted as Softmax; 3) we introduce our novel linear rectified router, which we denote with the abbreviation LR. The results are summarized in Table VI and visualized in Figure 8. As observed, while the Top-1 allocation MoE outperforms the standard LoRA (Var1), it is surpassed by the ”soft” allocation strategies (i.e.Var2 and Var3). This indicates the efficiency of employing an ensemble of experts to address a single task. Moreover, the comparison between the LR and Softmax routers (i.e. Var2 and Var3, Var4 and Var5) suggests that the LR routing strategy is superior. These results further underscore the effectiveness of our linear rectification routing strategy.
@@
Heatmap visualization of expert activation. In addition, we visualize the heatmaps of router outputs for both the Softmax and LR routing strategies across several middle transformer blocks. Figure 7 presents this comparative analysis, clearly demonstrating that our proposed Linear Rectified Router achieves a higher degree of sparsity in task allocation. Furthermore, we visualize the heatmaps of expert activation for the middle layer of top 40 transformer blocks across different tasks, as shown in Figure 10. The results indicate that different tasks engage different experts, with some tasks potentially sharing the same experts. These observations confirm that the RoDE model employs skill modules in a sparse manner and orchestrates skill sets in task-specific combinations, further substantiating the efficacy of our proposed RoDE approach.
Figure 10: Router Allocation Heatmaps for Different Tasks. The horizontal axis represents the router outputs, while the vertical axis represents the top 40 transformer blocks.
V-C3 Rank configuration of LoRA